OpenStudy (darkworld131):

The incidence of cystic fibrosis, a recessive genetic disorder in the Caucasian population of United States, is 1 in every 2,500 individuals. Find the number of heterozygous carriers. (p + q = 1, p2 + 2pq + q2 = 1) A:0.0392 B:0.0254 C:0.045

OpenStudy (anonymous):

This is not my strong point, but I would think that q2 = 1/2500, so q = sqrt(1/2500) and you can work it out from there.

OpenStudy (anonymous):

I THINK ITS C

OpenStudy (anonymous):

sorry if im wrong

OpenStudy (anonymous):

c is incorrect

OpenStudy (thelittlemerpmaid):

It is A